OpenStudy (george.g101):

Given that f(x) = 2x + 1 and g(x) = the quantity of 3x minus 1, divided by 2, solve for g(f(3)) A) 5 B) 7 C) 9 D) 10

OpenStudy (george.g101):

Please help @jigglypuff314 @agent0smith @YanaSidlinskiy

OpenStudy (yanasidlinskiy):

I helped you with a similar one yesterday. Try using that example and putting these numbers as I did yesterday

OpenStudy (george.g101):

I forgot how you laid it out lol, do you mind giving me the format and I'll plug in the numbers?

OpenStudy (yanasidlinskiy):

Hold on:)

OpenStudy (george.g101):

KK tyy :)

OpenStudy (yanasidlinskiy):

Refer to that: http://openstudy.com/users/george.g101#/updates/5762fbc2e4b02b7e8db80aed

OpenStudy (george.g101):

Ok so \[g(x)= 3(3)-1 \over 2\]

OpenStudy (george.g101):

g(x)=\[g(x)= 5/2\]

OpenStudy (george.g101):

Wait no 8/2= g(x)= 4

OpenStudy (george.g101):

SOS @agent0smith lol

OpenStudy (agent0smith):

Keep in mind you're finding g(f(3)); you need the INSIDE function first. Not the outside one.

OpenStudy (george.g101):

Sorry I'm really confused math is my weak spot =/

OpenStudy (agent0smith):

g( f(3) ) ^find f(3) first.

OpenStudy (george.g101):

f(3)= 2(3)+1? = f(3)= 7

OpenStudy (agent0smith):

Good now find g( f(3) )

OpenStudy (george.g101):

g(3)= 3(3)-1 over 2. = 8/2= g(3)=4?

OpenStudy (george.g101):

I'm stuck now :/

OpenStudy (agent0smith):

You're finding g( f(3) ) What was f(3)?

OpenStudy (george.g101):

7

OpenStudy (agent0smith):

So then replace f(3) with 7. You aren't finding g(3). You're finding g( f(3) )

OpenStudy (george.g101):

Wait I just worked it out the answer is '9' right?

OpenStudy (agent0smith):

Idk, show your work.

OpenStudy (george.g101):

Intelligence, I just summited it it I'll tell you if I was right. Thanks for your help:)

OpenStudy (agent0smith):

No... show your work.

OpenStudy (agent0smith):

I don't care if the answer is right or not. I'm saying show your work because if the work is right, the answer will be.

OpenStudy (agent0smith):

And on a test, teachers don't give full credit for just an answer. Work needs to be shown.